### v1

# Towards Global AI Inclusivity: A Large-Scale Multilingual Terminology Dataset

Abstract
The field of machine translation has achieved significant advancements, yet domain-specific terminology translation, particularly in AI, remains challenging. We introduced GIST, a large-scale multilingual AI terminology dataset containing 5K terms extracted from top AI conference papers spanning 2000 to 2023. The terms were translated into Arabic, Chinese, French, Japanese, and Russian using a hybrid framework that combines LLMs for extraction with human expertise for translation. The dataset’s quality was benchmarked against existing resources, demonstrating superior translation accuracy through crowdsourced evaluation. GIST was integrated into translation workflows using post-translation refinement methods that required no retraining, where LLM prompting consistently improved BLEU and COMET scores. A web demonstration on the ACL Anthology platform highlights its practical application, showcasing improved accessibility for non-English speakers. This work aims to address critical gaps in AI terminology resources and fosters global inclusivity and collaboration in AI research.
Towards Global AI Inclusivity:
A Large-Scale Multilingual Terminology Dataset
Jiarui Liu,1thanks: Equal contribution. Iman Ouzzani,1footnotemark: Wenkai Li,1footnotemark: Lechen Zhang,2 Tianyue Ou,1 Houda Bouamor,1 Zhijing Jin,3 Mona Diab1 1Carnegie Mellon University, 2University of Michigan, 3University of Toronto {jiaruil5, iouzzani, wenkail}@andrew.cmu.edu, zjin@cs.toronto.edu, mdiab@andrew.cmu.edu
1 Introduction
Figure 1: Direct translations of AI research papers and model cards using Google Translate generally offer fair quality but often fail to accurately translate AI-specific terminologies, potentially causing confusion or misunderstanding for readers. Our work addresses this issue by providing high-quality translations for a wide range of such terms, which can be efficiently integrated post-hoc to enhance the initial translations.
Figure 2: Overview of GIST creation. AI terminologies were extracted from awarded papers using an LLM and combined with existing terminology dictionaries. After data cleansing, translations into five languages were generated via crowdsourcing, and an LLM was used to select the best candidate translation. Dataset quality was evaluated against two baselines: majority-vote-selected translations and translations from the 60-60 evaluation set, demonstrating the superior quality of GIST.
The field of machine translation has made significant progress, with state-of-the-art models excelling across diverse tasks (Brown et al., 1990; Wu et al., 2016; Goyal et al., 2022; Haddow et al., 2022) and demonstrating effectiveness in translating between high-resource and low-resource languages (Yao and Wan, 2020; Costa-jussà et al., 2022; Ranathunga et al., 2023). Despite these successes, translating domain-specific scientific texts remains a persistent challenge, particularly for terminology translation (Cabré, 2010; Shuttleworth, 2014; Naveen and Trojovskỳ, 2024). General-purpose translation systems often falter in accurately translating terminologies, leading to loss of critical details (Dagan and Church, 1994; Haque et al., 2020a), or lead to misinterpretations (Chmutina et al., 2021; Yue et al., 2024).
In the field of AI, terms such as “Coreference Resolution” or “Explain-Away Effect,” are frequently translated incorrectly or inconsistently, undermining comprehension for global researchers and practitioners (Khuwaileh and Khwaileh, 2011; Tehseen et al., 2018), as illustrated in Figure 1. For non-English readers, who comprise a substantial portion of the global population and AI community (Ammon, 2003; Ding et al., 2023), such inaccuracies in AI research paper and blog translation hinder access to essential knowledge, stifling research innovation and collaboration across linguistic boundaries (Amano et al., 2023; Bahji et al., 2023). The implications extend to AI resources documentation, such as model and data cards hosted on platforms like Hugging Face,111huggingface.co/docs/hub/en/model-cards, huggingface.co/docs/hub/en/datasets-cards where errors in translated terminologies can lead to misunderstandings or incorrect usage of models and datasets. This highlights the urgent need for precise and standardized multilingual AI terminology resources to support equitable access to AI knowledge (Ahuja et al., 2023; Liu et al., 2024).
Existing efforts to address this gap have relied heavily on manual curation of terminologies by domain experts, a process that is time-consuming, resource-intensive, and difficult to scale (Wang et al., 2013; Freitag et al., 2021). Meanwhile, large language models (LLMs) have shown promise in automated terminology translation (Feng et al., 2024), while their outputs often misalign with human expert standards (Zhang et al., 2024), and different models often yield inconsistent translations (Banik et al., 2019; Prieto Ramos, 2021). The ACL 60-60 initiative in 2022 curated and translated AI-specific terms, and showcased potential for multilingual scientific communication and AI inclusivity (Salesky et al., 2023).2222022.aclweb.org/dispecialinitiative.html However, its small scale, narrow scope, and translation quality limitations underscore the need for a more comprehensive multilingual AI terminology resource.
To address this, we introduce Glossary of Multilingual AI Scientific Terminology (GIST), the first large-scale multilingual AI terminology dataset, compiling 5K AI-specific terms from award-winning papers at 18 top-tier AI conferences (2000–2023). Using a hybrid approach combining LLMs for extraction and verification with human expertise for translation, we provide high-quality translations into Arabic, Chinese, French, Japanese, and Russian (Abel and Meyer, 2013).
To enhance accessibility of our work, we integrate the curated terminologies into machine translation pipelines without requiring model retraining. We evaluate three post-translation refinement methods: prompting, word alignment and replacement, and constrained decoding. Our experiments reveal that the prompting method effectively integrates the curated dictionary, consistently improving translation quality as measured by BLEU and COMET scores. Additionally, we develop a website demonstration that incorporates our dataset into translations of ACL Anthology research papers, showcasing its practical applications for non-English speakers in the AI field.
The contributions of this work are summarized as follows:
  • We created GIST, the first large-scale multilingual AI terminology dataset, with 5K terms from 879 awarded top AI conference papers (2000–2023) and a broad domain coverage. [^3] Our data and code have been uploaded to the submission system, and will be open-sourced upon paper acceptance.
  • We developed an effective and efficient translation framework combining LLMs and human expertise to translate English terms into five languages, achieving high-quality results validated by automatic and human evaluations.
  • To enhance accessibility, we explored three approaches to integrate the curated terminologies into machine translation pipelines without retraining, and developed a website to showcase its practical use for AI paper translation.
By addressing the critical gap in multilingual AI terminology resources, this work contributes a robust solution to support equitable access to AI knowledge, fostering inclusivity and collaboration in the global research community.
2 Related Works
Multilingual Terminology Datasets
The ACL 60-60 evaluation sets represent an important effort in translating multilingual terminology from AI presentations, but are limited to just over 250 terms (Salesky et al., 2023).
Scientific terminology datasets are typically built using two approaches. The first approach relies on human multilingual experts for extracting and translating terminologies (Awadh, 2024; Kim et al., 2024). In cases of limited expert availability, human crowdsourcing with aggregation techniques has shown excellent results (Zaidan and Callison-Burch, 2011; Chan et al., 2023). The second approach uses ML and machine translation tools for automatic collection and translation, including log-likelihood comparisons (Haque et al., 2018), machine translation-based data synthesis (Haque et al., 2020b; Fernando et al., 2020; Manzini et al., 2022; Moslem et al., 2023), platform-based terminology linkers (Arcan et al., 2014), classifier training (Jin et al., 2013; Schumann and Martínez Alonso, 2018), and prompting LLMs (Nishio et al., 2024; Shamsabadi et al., 2024).
However, fully human-driven approaches can be costly for large-scale multilingual datasets, while fully automated ones often lack accuracy (Giguere, 2023). To address this, our framework integrates LLM-based extraction, human filtering, human translation, LLM validation, and merging with existing dictionaries, efficiently and effectively expanding existing terminology datasets.
Integrating Domain Terminology into Machine Translation
Integrating newly collected domain terminology into machine translation systems has led to a variety of research efforts. One common approach involves training methods, such as augmenting training data with input-output pairs that include terminologies for supervised fine-tuning (Dinu et al., 2019; Niehues, 2021), or modifying model architectures to enhance terminology awareness (Dinu et al., 2019; Conia et al., 2024). However, these training-based approaches are inefficient for adapting to new terminologies, limiting their real-world applicability.
To tackle this, terminology-aware decoding methods have emerged as a more flexible alternative. These methods, which include variants of constrained beam search (Anderson et al., 2016; Hokamp and Liu, 2017; Chatterjee et al., 2017; Hasler et al., 2018; Post and Vilar, 2018), slightly compromise translation accuracy for greater adaptability. Another strategy is post-hoc editing of generated translations, which typically employs word alignment techniques to identify and replace term translations in the output (Zenkel et al., 2019; Chen et al., 2020; Ferrando et al., 2022). Recently, LLMs have been leveraged to integrate expected terminology directly into the translations (Bogoychev and Chen, 2023). Our work explores multiple terminology integration approaches, including constrained beam search, decoding logits adjustment, word alignment and replacement, and refinement through prompting.
3 GIST Dataset Construction
We constructed GIST, a dataset comprising around 5K English AI terminologies and their translations into Arabic, Chinese, French, Japanese, and Russian. The basic lexical statistics of the dataset are presented in Table 1. We detail our terminology collection process in Section 3.1 and the translation procedure in Section 3.2.
Arabic Chinese French Japanese Russian
# Terms 4,844 6,426 6,527 4,770 5,167
Unique En Words 2,470 3,244 3,470 2,424 2,615
Unique Tgt Words 3,161 2,838 4,036 2,050 4,210
En Words/Term 2.02 2.05 2.07 2.02 2.01
Tgt Words/Term 2.36 2.26 2.68 2.53 2.16
En Chars/Term 16.99 17.26 17.44 16.96 16.94
Tgt Chars/Term 15.22 4.66 21.27 6.89 20.20
Table 1: Lexical statistics of the dataset across languages. “En” denotes English, and “Tgt” denotes the target language. Terms are tokenized into words using nltk.word_tokenize for English, Arabic, French, and Russian, jieba for Chinese, and MeCab Owakati for Japanese. Statistics with standard deviations are presented in Table 3.
3.1 Terminology Collection
Our dataset follows the 60-60 initiative that aims to collect scientific terminologies in the AI field. We source AI terminology from two primary channels: AI research papers published online and existing AI terminology dictionaries. For the research papers, our objective is to compile a substantial number of terms from high-quality AI papers spanning a long time frame. To identify representative papers, a natural approach would be to crawl the most cited papers using online search engines or platforms. However, no tools or APIs currently enable this. Instead, we focused on awarded papers announced on the websites of top AI conferences. Specifically, we collected all awarded papers, spanning awards such as Best Paper, Outstanding Paper, and other recognitions from the venues listed in Table 4, and covering the years 2000 to 2023. This approach ensures comprehensive coverage of recognized research work. In total, we collected 879 paper PDFs from arXiv and other online repositories and processed them into text files using SciPDF. [^4] https://github.com/titipata/scipdf_parser As we analyze later in Section 5, this collection strategy provides broad terminology and domain coverage of AI terminologies.
As supported by previous research (Xu et al., 2024; Dagdelen et al., 2024; Shamsabadi et al., 2024), recent LLMs such as LLaMA 3 (Dubey et al., 2024) have demonstrated strong capabilities in scientific terminology extraction tasks. While the definition of “AI” lacks a clear and universally agreed boundary, LLMs are trained on vast datasets that reflect human knowledge, enabling them to classify terms as AI-related based on their contextual relevance. Leveraging this, we utilized LLaMA-3-70B-Instruct to extract AI terminologies from award-winning papers, providing specific instructions to guide the extraction process.
To define AI-specific terminology, we imposed the following criteria: (1) the term must be a noun or noun phrase, (2) it should be specialized to AI, encompassing core concepts, methods, models, algorithms, or systems, and (3) it should have either no meaning or a distinct meaning outside the AI domain. We processed the text in sentence chunks of up to 64 words to stay within LLaMA’s optimal context-handling capabilities. Additionally, for each unique term, we recorded up to three different contexts during extraction to ensure sufficient contextual diversity.
After extraction, we performed multiple quality assurance steps. We removed terms that appeared in only one paper to ensure representativeness, excluded abbreviations and terms starting with special characters, and filtered out non-noun phrases. Duplicates were eliminated, and GPT-4o (Hurst et al., 2024) was employed to further refine the list by filtering out non-AI terms based on the same criteria. Finally, three AI domain experts reviewed the terms to remove any remaining unqualified entries. To enhance the dataset, we also integrated terminologies from external sources, including the 60-60 initiative dataset, government websites, Wikipedia, and other online resources. Full details of the terminology collection process are provided in Section A.2.
3.2 Terminology Translation
To evaluate whether state-of-the-art LLMs can effectively perform terminology translation as a generation task, we initially experimented with two advanced LLMs and one API available at the time, including Claude 3 Sonnet (Anthropic, 2024), GPT-3.5-Turbo (OpenAI, 2023), and Google Translate API (Wu et al., 2016). The procedure and the prompt used for these translations is presented in Section A.3. We measured the agreement among the three methods using exact match. However, as shown in Table 5, the three-model agreement ratio was only around 15% for most languages except Chinese, and the two-model agreement ratio was only about 40%. These results reveal significant inconsistencies in the translations produced by these systems, and highlight the need for human inputs to ensure reliable translations.
Given these findings, we opted for human annotation to ensure translation accuracy. To achieve this, we utilized Amazon Mechanical Turk (MTurk) for crowdsourced annotations. A demonstration of the MTurk task is shown in Figure 6. We instructed participants to take on the task only if they specialized in AI and were fluent in both English and one of the target languages. Annotators were tasked with generating accurate translations for each AI terminology, with relevant contexts provided and the terminology highlighted in yellow in contexts. To maintain quality, we implemented a rigorous qualification process. Annotators were first tested on a toy set of 10 carefully selected AI terms, and only those who performed well were allowed to proceed with the full task. Additionally, we monitored submissions daily and filtered out participants who provided random or low-quality translations during the annotation process. For each term, we collected 10 annotations per target language. The total cost is around \$1000.
Finally, we used GPT-4o to select the best translation from the annotators’ submissions for each term, ensuring high-quality results for our final dataset. As analyzed in Section 4.1, leveraging GPT-4o was crucial for maintaining the quality of the translations.
4 Dataset Quality Assessments
To thoroughly evaluate the translation quality in GIST, we conducted two additional crowdsourced rating tasks: (1) In Section 4.1, we investigated whether using GPT-4o was necessary to select the best translation candidate; (2) In Section 4.2, we compared the quality of our dataset with the evaluation set from the 60-60 initiative.
4.1 Task 1: Is Our LLM + Human Framework Necessary?
To ensure the selection of the best candidate from the annotators’ generations, we explored two methods for candidate selection. The first method only uses GPT-4o to select the best translation candidate among 10 annotations for all terms. The second method relies on majority voting among the 10 annotators: a translation candidate is selected if it appears in more than 5 out of 10 annotations. In cases where no majority is reached, GPT-4o is prompted to select the best translation candidate.
To evaluate these approaches, we conducted a separate MTurk task, involving a different group of participants, to compare the two methods on a randomly sampled subset of 200 terms per language. Participants were asked to choose one of four options: (A) Both translations are good; (B) Method 1 translation is better; (C) Method 2 translation is better; (D) Both translations are bad. For each language, we collected 5 annotations per term. As shown in Table 6, GPT-4o’s candidate selection consistently outperformed majority voting across all five languages. These results highlight the necessity of GPT-4o in achieving high-quality translations for crowdsourced annotations.
However, human involvement remains indispensable. As discussed in Section 3.2, state-of-the-art LLMs and machine translation systems fail to provide consistent answers, rendering automatic translation through agreement across multiple models infeasible. Thus, by combining the expertise of LLMs as verifiers with humans as input sources, we achieve efficient, accurate, and reliable translations.
4.2 Task 2: Is Our Dataset Translation Better than 60-60?
The 60-60 initiative dataset also focuses on AI terminology translations and includes the five languages we considered. This overlap allows us to intersect the English terms in our dataset with those in the 60-60 evaluation dataset and compare their respective translations. We retrieved the translations from both datasets and conducted a comparative assessment through a crowdsourced evaluation, same as the process detailed in Section 4.1.
The results presented in Table 7 show that the translations in our dataset consistently and significantly outperform those from the 60-60 dataset across all five languages. Additionally, as shown in Table 8, annotators demonstrated fair agreement when rating the Arabic translations, and moderate agreement, with Fleiss’ Kappa values ranging from 0.4 to 0.5, for translations in the other four languages. These findings underscore the superior quality of translations in our dataset compared to the 60-60 dataset.
5 Dataset Coverage Assessments
Domain Coverage
We first investigate the AI domains covered by the collected terms in GIST and their distributions. We use GPT-4o-mini (Hurst et al., 2024) to identify the specific AI domains for each term, following the taxonomy proposed by Ding et al. (2023), which clusters the research domains of AI scholars. Figure 8 shows the distribution of the top six AI domains in GIST, with terms most frequently categorized into statistics, mathematics, computer science (CS), natural language processing (NLP), data science (DS), and computer vision (CV). We further embed all AI terminologies in GIST using the all-MiniLM-L6-v2 model (Reimers and Gurevych, 2019) and apply Uniform Manifold Approximation and Projection (UMAP) (McInnes et al., 2018) for dimensionality reduction to a two-dimensional space. The visualization in Figure 9 reveals that terms in domains such as NLP, CV, statistics, and mathematics form distinct clusters, while terms in CS and DS are more dispersed.
Terminology Coverage
We claim that our awarded paper dataset provides essential coverage of the terms in our terminology dictionary, where many terms are actually repeatedly extracted multiple times from different awarded papers. We present the rarefaction curve in Figure 10, which visually demonstrates the relationship between the number of extracted terms and the dataset size.
To statistically validate our claim, we conducted a one-sided one-sample t-test to compare the mean coverage ratio of the terminology dictionary. Specifically, we tested whether the mean coverage ratio of randomly sampled subsets, constituting 60% of the terminology dictionary, is significantly above 80%. The null hypothesis ($H_{0}$) assumes that the mean coverage ratio of these subsets is less than or equal to 80%, while the alternative hypothesis ($H_{a}$) assumes that the mean coverage ratio exceeds 80%. Using 1,000 random samples, the analysis yielded a t-statistic of 64.78 and a p-value of 0, rejecting the null hypothesis. These results provide statistical evidence supporting the sufficiency of our dataset in covering the terminology dictionary.
Eval Set Model Metric Arabic Chinese French Japanese Russian
60-60 aya-expanse-8B BLEU 20.11 + 1.23 + 0.18 27.31 + 1.33 + 0.24 33.05 + 2.46 + 0.20 14.59 + 0.61 + 0.32 16.59 + 1.59 - 0.05
COMET 81.96 + 0.71 - 0.52 83.43 + 1.57 + 0.08 81.83 + 1.06 - 0.11 88.54 + 0.32 - 0.01 82.27 + 0.69 - 2.02
aya-23-8B BLEU 19.98 + 0.54 - 0.21 26.08 + 0.47 + 0.39 33.85 + 2.28 - 0.11 15.06 + 0.87 + 0.36 15.77 + 1.05 + 0.37
COMET 84.02 + 0.81 - 0.24 85.12 + 0.58 + 0.38 82.40 + 0.94 - 0.15 87.92 + 0.50 + 0.09 81.91 + 0.40 - 2.26
gpt-4o-mini BLEU 23.58 + 1.07 - 0.00 32.64 + 1.60 + 0.66 40.80 + 3.08 + 0.50 21.46 + 0.64 + 0.19 17.25 + 1.07 - 0.13
COMET 85.77 + 0.69 - 0.44 87.30 + 0.48 + 0.26 84.56 + 0.68 - 0.04 89.96 + 0.14 + 0.01 83.68 + 0.38 - 2.29
nllb BLEU 22.38 + 1.37 + 0.64 17.29 + 1.92 + 1.02 34.93 + 2.86 + 0.21 6.19 + 2.42 + 0.53 17.30 + 1.54 + 1.07
COMET 83.52 + 0.83 - 0.45 78.22 + 2.95 + 0.73 82.83 + 1.00 - 0.19 77.82 + 3.80 + 0.39 81.41 + 0.97 - 1.54
seamless BLEU 23.13 + 1.16 - 0.03 26.26 + 0.97 + 0.80 40.04 + 2.08 - 0.57 14.56 + 0.74 + 0.05 17.18 + 1.71 + 1.17
COMET 84.07 + 0.94 - 0.38 83.44 + 1.48 + 0.50 83.86 + 0.78 - 0.07 85.05 + 1.06 + 0.16 82.33 + 0.56 - 1.87
AI Papers & Model Cards aya-expanse-8B BLEU 11.47 + 0.37 + 0.10 12.04 + 0.94 + 0.17 18.84 + 1.71 - 0.85 8.11 - 0.03 + 0.04 13.84 + 0.21 + 0.32
COMET 80.98 + 0.46 - 0.51 82.42 + 0.56 - 0.01 81.16 + 0.36 - 0.79 85.48 + 0.34 + 0.07 82.76 + 0.47 - 2.10
aya-23-8B BLEU 14.28 + 0.81 + 0.28 14.50 + 0.39 + 0.19 24.49 + 2.36 + 0.08 9.22 + 0.23 + 0.35 16.39 + 1.36 + 0.61
COMET 81.55 + 1.03 - 0.62 83.88 + 0.68 + 0.04 82.55 + 1.22 - 0.70 84.42 + 0.81 + 0.02 82.72 + 1.14 - 2.08
gpt-4o-mini BLEU 14.37 + 0.53 - 0.42 17.21 + 1.22 + 1.39 24.45 + 4.38 + 1.28 10.55 + 0.05 + 0.03 18.02 + 1.52 + 0.40
COMET 83.56 + 0.86 - 0.19 86.08 + 0.25 - 0.13 84.75 + 0.33 - 1.03 87.91 + 0.16 - 0.01 84.92 + 0.44 - 2.02
nllb BLEU 15.42 - 0.31 - 0.77 10.24 + 2.19 + 2.07 22.68 + 2.73 + 0.90 8.24 + 1.08 + 0.88 19.18 - 0.10 - 0.24
COMET 81.23 + 1.28 - 0.50 80.19 + 1.61 + 0.38 78.70 + 3.81 - 1.59 83.05 + 1.71 + 0.70 80.46 + 2.80 - 2.38
seamless BLEU 15.38 + 1.09 + 0.45 13.67 + 1.10 + 0.73 24.34 + 5.21 + 1.49 9.42 + 0.56 + 0.42 18.43 + 0.95 + 0.35
COMET 81.96 + 1.18 - 0.39 80.70 + 2.18 + 0.16 83.76 + 0.97 - 0.94 83.70 + 0.88 + 0.10 83.12 + 1.50 - 1.79
Table 2: Experiment results on two evaluation sets across five models and five languages using BLEU and COMET. The first value represents the scores for direct translation using the models. The second value indicates the improvement or decline when applying the prompting-powered refinement method compared to direct translation. The third value indicates the improvement or decline when using the word alignment method compared to direct translation.
6 Experiments
We explore terminology integration approaches that do not need model retraining. We quantitatively evaluate two methods: terminology refinement via post-hoc LLM prompting and terminology substitution guided by word alignment. Additionally, we conduct a qualitative evaluation of traditional decoding techniques, including constrained beam search and token-level logits adjustment.
6.1 Methods
Terminology Refinement: LLM Prompting
With the recent advancements in multilingual and instruction-following capabilities of LLMs, it is now possible to leverage these models to refine translations. Specifically, we prompt the LLMs to revise the initial translation produced by a machine translation model, incorporating relevant term translations from our terminology dictionary within the provided context. For this refinement, we employ GPT-4o-mini and use the prompt illustrated in Figure 11.
Terminology Substitution: Word Alignment
We apply the word alignment approach introduced by Dou and Neubig (2021) to identify and substitute term translations within the output. We use the multilingual BERT base model (Devlin et al., 2018) to to tokenize both source and target sentences, and then processes these inputs through its hidden layers to produce contextualized embeddings. We determine alignments by computing dot-product similarities between source and target token embeddings. High-confidence alignments are then filtered using a threshold of 1e-4, and subword alignments are aggregated to generate word-level mappings. Lastly, we directly replace identified term translations with those from GIST.
Terminology-Aware Decoding Approaches
For constrained beam search, we use the Hugging Face implementation, which enforces the inclusion of term translations from GIST in the output. [^5] https://huggingface.co/blog/constrained-beam-search For token-level adjustment, we first identify AI terminologies and their corresponding expected translations from GIST. We then modify the output logits of these tokens by increasing them by factors of 10/7, 10/8, and 10/9 to prioritize their selection during decoding.
6.2 Experiment Setup
Evaluation Set
We conducted experiments on two evaluation sets: the 60-60 Evaluation Set and the AI Papers and Model Cards Evaluation Set. The first evaluation set is from the 60-60 initiative, providing terminology translations into five target languages. Based on our analysis in Section 4.2, which indicates that our dataset offers higher quality than the 60-60 set, we enhanced their labels by replacing AI terminologies in their translations with GPT-4o while maintaining grammatical correctness and ensuring no loss in translation quality. This process generated updated ground truth labels for the evaluation.
The second evaluation set was manually created by combining text from two sources: 50 held-out AI research papers and 50 model cards generated by (Liu et al., 2024). From this set, we randomly sampled 500 English text chunks from it for evaluation. To create ground truth labels for this set, we used Google Translate, a state-of-the-art machine translation model (Zhu et al., 2023; Santosa et al., 2024), to produce initial translations. We then prompted GPT-4o to refine these translations by updating AI terminology. This evaluation set was designed to explore the application of our AI terminology dictionary in two major domains: AI research papers and model cards.
Models
We evaluated the following models: gpt-4o-mini (Hurst et al., 2024), hf-seamless-m4t-large (Barrault et al., 2023), nllb-200-3.3B (Costa-jussà et al., 2022), aya-23-8B (Aryabumi et al., 2024), and aya-expanse-8B (Dang et al., 2024). While the prompting method applies to all models, the word alignment and decoding methods are not applicable to gpt-4o-mini, as it does not provide access to its model weights.
Evaluation Metrics
We adopt BLEU (Papineni et al., 2002) and COMET metrics (Rei et al., 2020) for the quantitative evaluation for translation quality, following the methodology of Salesky et al. (2023). We utilize the wmt22-comet-da model (Rei et al., 2022).
6.3 Experiment Results
Figure 3: The interface demonstrates a paper example from (Tang et al., 2024) and introduces a new translation feature on the ACL Anthology website. Non-English speakers can click the “Translate” button at the bottom right of the webpage and select their preferred language for translation. The page dynamically displays two translations side by side: one using direct translation from a machine translation model (e.g., seamless in the figure) and the other enhanced with prompting-powered refinement applied to the left translation. Highlighted text indicates updated terminologies integrated from GIST, showcasing improved translation performance.
Table 2 presents the quantitative evaluation results under the experimental settings described earlier. We draw several key observations:
First, the results across the two evaluation sets show consistency, indicating the robustness of the findings. Among the models, gpt-4o-mini achieves the best overall translation performance. Among the remaining four models, differences in performance are less pronounced.
Several trends emerge when comparing different terminology integration approaches. The prompting-powered refinement method consistently outperforms direct translation across nearly all languages, models, and evaluation metrics, which highlights its effectiveness in effectively incorporating AI-specific terminology into translations.
The word alignment method demonstrates mixed performance, showing improvements for Chinese and Japanese translations but leading to declines for Arabic, French, and Russian. This discrepancy is likely due to linguistic differences: Chinese and Japanese exhibit minimal morphological changes, allowing straightforward substitution of terminology without disrupting surrounding syntax; In contrast, languages like Arabic often require agreement in gender, number, and syntactic roles, making noun replacement more challenging and prone to errors.
We also conducted a qualitative manual evaluation of the generation results produced by constrained beam search and logit adjustment methods. Unfortunately, both approaches exhibited extremely slow performance, running approximately 100 times slower than post-hoc methods. Furthermore, forcing specific word ids to appear in the output did not perform well, suffering from similar issues as the word alignment method in disrupting syntactical dependencies. During manual inspection, we observed that constrained beam search behaved similarly to the logit adjustment method with a large scaling factor: the generated sentences often failed to maintain proper grammar, and terminologies were frequently repeated multiple times within the same output. Conversely, with a small scaling factor in the logit adjustment method, the terms were often omitted entirely, demonstrating little to no effect on the output. These findings highlight the limitations in these traditional decoding approaches compared to the prompting-powered refinement method, as well as the importance of tailoring terminology integration approaches to the linguistic characteristics of target languages.
7 Website Demonstration
To facilitate real-world usage of our AI terminology dictionary, we built on the 60-60 initiative by modifying the ACL Anthology website layout and introducing a new terminology translation feature, [^6] https://acl6060.org/ as illustrated in Figure 3. The ACL Anthology website was chosen for this demonstration due to its extensive collection of AI-related research papers, making it an ideal platform to showcase the potential impact of our work. In this demonstration, translations of key terms are refined and standardized based on the terminology dictionary we developed, providing more accurate terminology translations. This enhancement represents a step toward improving access to AI knowledge for non-English speakers by offering a consistent and reliable translation system, and broadens the accessibility of AI research to a global audience.
8 Conclusion
We present GIST, a large-scale multilingual AI terminology dataset addressing gaps in translating AI-specific terms. Combining LLM-based extraction and validation with human expertise, it includes 5K English terms with translations into Arabic, Chinese, French, Japanese, and Russian, surpassing the ACL 60-60 benchmark. LLM prompting proved effective for post-hoc terminology integration, improving translation quality across metrics like BLEU and COMET. We also provide a website demonstration to enhance the accessibility of our work for non-English speakers, supporting equitable AI knowledge access and fostering global collaboration in AI research.
Limitations
This work is subject to several limitations. First, our dataset assumes a one-to-one correspondence between English terms and their translations, which does not account for cases where multiple equally valid translations exist for a single term. This simplification may overlook the nuanced variations in AI terminology usage across languages.
Despite significant efforts in collecting terminologies, our coverage is not exhaustive. The field of artificial intelligence lacks a well-defined boundary, making it challenging to ensure comprehensive inclusion of all relevant AI terms. Additionally, while we provide translations for five widely used languages, this represents only a subset of the global linguistic diversity and leaves many other languages unaddressed.
The evaluation set was created using Google Translate and GPT-4, tools whose translation accuracy has not been verified in our work by human experts. This reliance might introduce potential errors in the evaluation data, affecting its reliability.
Nonetheless, we hope this work inspires the community to further advance the creation and refinement of multilingual AI terminology dictionaries, addressing these limitations and extending coverage to more languages and domains.
Ethical Considerations
We manually collected awarded papers from official conference websites. Only papers available under open-source licenses for research use were downloaded. Similarly, model card contents used in this study were sourced from openly shared materials by their respective authors.
During data processing, we ensured that no personal information, such as human names irrelevant to AI methods or metrics, was included in the dataset. Automatic and manual reviews were conducted to verify the exclusion of any sensitive or private details.
For term translation, we employed LLMs to assist with extraction and validation. While we acknowledge that LLMs may exhibit biases when selecting the best translations, manual evaluations confirmed the superior performance of LLMs in this task.
In crowdsourced experiments, we respected participant privacy by not collecting any demographic information. All contributors were fairly compensated, with the total cost for annotations amounting to approximately \$1,000. Additionally, our collected terminology dataset does not involve ethically sensitive or controversial content, focusing exclusively on technical terms relevant to AI.
Appendix A Additional Dataset Details
A.1 Additional Dataset Statistics
Arabic Chinese French Japanese Russian
# Terms 4844 6426 6527 4770 5167
Unique En Words 2470 3244 3470 2424 2615
Unique Tgt Words 3161 2838 4036 2050 4210
En Words/Term $2.02\pm 0.59$ $2.05\pm 0.68$ $2.07\pm 0.67$ $2.02\pm 0.58$ $2.01\pm 0.59$
Tgt Words/Term $2.36\pm 0.83$ $2.26\pm 0.90$ $2.68\pm 1.19$ $2.53\pm 0.98$ $2.16\pm 0.80$
En Chars/Term $16.99\pm 5.97$ $17.26\pm 6.60$ $17.44\pm 6.57$ $16.96\pm 5.91$ $16.94\pm 5.90$
Tgt Chars/Term $15.22\pm 5.66$ $4.66\pm 1.96$ $21.27\pm 8.49$ $6.89\pm 3.16$ $20.20\pm 7.83$
Table 3: Lexical statistics of the dataset across languages, including standard deviations. “En” denotes English, and “Tgt” denotes the target language. Terms are tokenized into words using nltk.word_tokenize for English, Arabic, French, and Russian, jieba for Chinese, and the MeCab Owakati tokenizer for Japanese.
To better understand the composition and characteristics of the dataset, we performed a comprehensive statistical analysis across multiple dimensions, including domain distribution, semantic clustering, and lexical structure across five languages: Arabic, Chinese, French, Japanese, and Russian. Below, we present key findings through detailed visualizations and lexical statistics.
Lexical Statistics.
Table 3 summarizes the lexical characteristics of the dataset across the five languages. Key metrics include:
  • Number of Terms: The dataset contains 4,844 to 6,527 terms per language, with French and Chinese having the largest repositories.
  • Unique Words: English terms comprise 2,400–3,400 unique words across datasets. Target languages exhibit varying lexical diversity, with French and Russian showing higher uniqueness due to linguistic richness.
  • Words per Term: On average, English terms consist of approximately two words ($\sim 2.02$), while target languages show higher variability. French terms, for instance, require more words ($2.68\pm 1.19$), reflecting language-specific expansion during translation.
  • Characters per Term: English terms maintain consistent lengths ($\sim 17$ characters), while target languages vary significantly. For example, Chinese terms are concise ($4.66\pm 1.96$) due to its logographic script, whereas French ($21.27\pm 8.49$) and Russian ($20.20\pm 7.83$) terms are longer, reflecting the morphology of these languages.
To tokenize terms, we utilized nltk.word_tokenize for English, Arabic, French, and Russian; jieba for Chinese; and MeCab Owakati for Japanese. These tools ensured language-specific tokenization accuracy, enabling detailed lexical analysis.
Conclusion.
The statistical analysis highlights the diversity and interdisciplinary nature of the dataset. Figures 8 and 9 illustrate domain-wise distributions and semantic clusters, while Table 3 quantifies lexical variations across languages. Together, these findings provide a robust understanding of the dataset’s structure, supporting its utility for multilingual and domain-specific AI applications.
A.2 Terminology Collection Details
We integrate terms from the Wiki-Pedia Glossary of AI777en.wikipedia.org/wiki/Glossary_of_artificial_intelligence, which serves as a comprehensive starting point for artificial intelligence (AI)-related terminology, ensuring alignment with globally recognized concepts and definitions. To enhance multilingual coverage and domain relevance, we also include terms from several other specialized AI terminology initiatives across different languages:
Arabic AI Dictionary888ai.gov.ae/ar/ai-dictionary/: Published by the Arabic Government AI Office, this resource aims to elevate the status of the Arabic language in AI, standardize terminology, reduce linguistic ambiguity, and foster better integration of Arabic speakers into the global AI community. By providing accurate translations and clear definitions for English AI terms, this dictionary promotes knowledge dissemination and encourages collaboration within the Arabic-speaking AI ecosystem.
Chinese GitBook AI Term Database [^9] https://jiqizhixin.gitbook.io/artificial-intelligence-terminology-database: Developed by Jiqizhixin (Machine Heart), this database represents an extensive effort to document technical terms encountered during the translation of AI articles and research papers. Starting with practical usage, the project has evolved to incorporate domain-specific expansions based on authoritative textbooks and expert input, offering the Chinese AI community a unified and precise reference for both academic and industrial applications.
French AI Dictionary [^10] https://datafranca.org/wiki/Cat%C3%A9gorie:
GRAND_LEXIQUE_FRAN%C3%87AIS
The first comprehensive French reference tool for data science and AI, this dictionary addresses the needs of public service, commerce, research, and education. It aims to bridge the gap between French and English AI terminologies, ensuring accessibility and standardization for French-speaking professionals and researchers.
Russian AI Dictionary [^11] https://www.raai.org/pages/UGFnZVR5cGU6MTAwMw==: This initiative captures the interdisciplinary nature of AI by including terminology drawn from fields such as logic, psychology, linguistics, and cybernetics. Leveraging contributions from Russian and Soviet experts, this dictionary emphasizes the frequency and relevance of terms within AI-specific contexts, providing a culturally adapted yet globally aligned resource for Russian-speaking researchers.
Figure 4: Prompts for extracting AI terminologies with LLaMA-3-70B-Instruct.
A.3 Terminology Translation Details
Terminology Dataset Collection
In this section, we provide a detailed description of our methodology for creating the multilingual AI terminology dictionary and the associated translation experiments. The terms were extracted from papers published in the top AI conferences as shown in Table 4 across various fields including Artificial Intelligence, Computer Vision, Machine Learning, Natural Language Processing, and Web & Information Retrieval. These conferences represent the leading venues in their respective domains.
Category Selected Conferences
AI AAAI, IJCAI
CV CVPR, ECCV, ICCV
ML ICLR, ICML, NeurIPS, KDD
NLP ACL, EMNLP, NAACL, EACL
LREC, COLING, CoNLL
Web & IR SIGIR, WWW
Table 4: Top AI conferences included in our dataset collection.
Prompt Design for Translation
To ensure consistent and high-quality translations, we used carefully designed prompts for GPT-3.5-Turbo and Claude 3 Sonnet. We provided sentence contexts relevant to each AI terminology, split into a maximum of three chunks, each containing up to 64 words. As shown in Figure 5, the translation prompt asks the model to translate an AI-specific term into a target language, with context provided to clarify the meaning. If the term is an abbreviation or a technical term that should remain in its original form to avoid confusion, the models are instructed to retain the English term.
Figure 5: Prompts for translating AI terminologies with GPT-3.5-Turbo and Claude 3 Sonnet.
Arabic Chinese French Japanese Russian
Three-Model Agreement Ratio 10.11% 42.71% 9.86% 16.60% 17.93%
Two-Model Agreement Ratio 30.52% 36.82% 45.15% 40.44% 38.36%
Table 5: Translation agreement ratios among three models (Claude 3 Sonnet, Google Translate API, and GPT-3.5-Turbo) for five target languages. The table shows the ratio of terms where all three models agree (Three-Model Agreement Ratio) and the ratio where any two models agree (Two-Model Agreement Ratio).
Translation Agreement Analysis
We evaluated the consistency of translations across three models: Claude 3 Sonnet, GPT-3.5-Turbo, and Google Translate API. Table 5 summarizes the agreement ratios for five target languages: Arabic, Chinese, French, Japanese, and Russian. The results indicate significant variation in agreement ratios across languages, with Chinese achieving the highest three-model agreement (42.71%) and Arabic having the lowest (10.11%). The detailed discussion is in Section 3.2.
Human Translation via Mturk
To further validate the translations, we employed human translator through the MTurk platform. Each terminology was translated by 10 independent annotators, resulting in a set of 10 translation candidates for each term. The platform layout and guidance for helping translators give these translations is shown in Figure 6.
Figure 6: The MTurk layout demonstration for AI terminology translation generation task.
Figure 7: The MTurk layout demonstration for evaluating AI terminology translations. The layout is used for two tasks: (1) comparing the GPT-4o-selected candidate with the majority-voted candidate, and (2) comparing the translations in our dataset with those in the 60-60 initiative evaluation set.
A.4 Dataset Quality Assessment Details
To evaluate the quality of translations in our dataset, we conducted two tasks involving human annotators. Annotators were presented with pairs of translations generated by different methods and were tasked with evaluating their relative quality. The analysis of their ratings is summarized in Tables 6, 7, and 8. Below, we describe the findings in detail.
Task 1: Comparison Between GPT-4o-Selected Candidates and Majority-Voted Translations
In Task 1, annotators compared translations selected by GPT-4o (Method 1) with majority-voted translations generated by human annotators (Method 2). The choices included:
A. Both translations are good B. Method 1 translation is better C. Method 2 translation is better D. Both translations are bad
Arabic 739, 45.76% 461, 28.54% 329, 20.37% 72, 4.46%
Chinese 468, 50.59% 266, 28.76% 163, 17.62% 25, 2.70%
French 438, 48.67% 274, 30.44% 170, 18.89% 17, 1.89%
Japanese 587, 56.99% 251, 24.37% 159, 15.44% 27, 2.62%
Russian 626, 54.43% 348, 30.26% 150, 13.04% 14, 1.22%
Table 6: Distribution of annotators’ choices for AI terminology ratings in Task 1, comparing the GPT-4o-selected candidate (Method 1) with the majority-voted candidate (Method 2). The table reports the total counts and corresponding ratios for each choice.
Table 6 summarizes the distribution of annotators’ choices for Task 1. Across all languages, the majority of annotations fell into category A, where both translations were rated as good. Japanese showed the highest percentage of agreement in this category (56.99%), followed by Russian (54.43%) and Chinese (50.59%). Cases where Method 1 translations were rated as better (category B) ranged from 24.37% (Japanese) to 30.26% (Russian). Similarly, cases where Method 2 translations were rated as better (category C) ranged from 13.04% (Russian) to 20.37% (Arabic). The lowest percentage of responses was observed in category D, where both translations were rated as bad, accounting for less than 5% of responses across all languages.
A. Both translations are good B. Method 1 translation is better C. Method 2 translation is better D. Both translations are bad
Arabic 376, 46.42% 238, 29.38% 143, 17.65% 46, 5.68%
Chinese 197, 37.17% 228, 43.02% 85, 16.04% 17, 3.21%
French 152, 39.48% 168, 43.64% 53, 13.77% 10, 2.60%
Japanese 295, 57.28% 162, 31.46% 36, 6.99% 21, 4.08%
Russian 172, 39.09% 198, 45.00% 38, 8.64% 25, 5.68%
Table 7: Distribution of annotators’ choices for AI terminology ratings in Task 2, comparing the translations in our dataset (Method 1) with those in the 60-60 initiative evaluation set (Method 2). The table reports the total counts and corresponding ratios for each choice.
Task 2: Comparison Between Dataset Translations and 60-60 Initiative Translations
In Task 2, annotators compared translations from our dataset (Method 1) with those in the 60-60 initiative evaluation set (Method 2). The distribution of choices is shown in Table 7. In this task, a higher percentage of annotators preferred translations from Method 2 (category B) for most languages, especially in Chinese (43.02%), French (43.64%), and Russian (45.00%). However, Japanese translations from Method 1 had a significantly higher percentage in category A, with 57.28% of annotators agreeing that both translations were good. Instances where both translations were rated as bad (category D) remained low across all languages, ranging from 2.60% to 5.68%.
Arabic Chinese French Japanese Russian
Task 1 0.30 0.22 0.37 0.21 0.39
Task 2 0.22 0.41 0.50 0.39 0.41
Table 8: Fleiss’ Kappa scores for inter-annotator agreement among 5 annotators across each question in Task 1 and Task 2.
Inter-Annotator Agreement
To measure inter-annotator agreement, we calculated Fleiss’ Kappa scores for each language in both tasks. Table 8 reports these values. Fleiss’ Kappa values between $0.400.20\leq\kappa<$ 0.40 indicate fair agreement, while values between $0.600.40\leq\kappa<$ 0.60 indicate moderate agreement. In Task 1, Kappa scores ranged from 0.21 (Japanese) to 0.39 (Russian), showing fair agreement across languages. In Task 2, Kappa scores improved, ranging from 0.39 (Japanese) to 0.50 (French), indicating moderate agreement for most languages.
Conclusion
The results demonstrate that our dataset’s translations are of high quality, with a majority of annotators rating them as good. While inter-annotator agreement was fair in Task 1, moderate agreement was observed in Task 2, highlighting the robustness of our dataset compared to established benchmarks.
A.5 Dataset Coverage Assessment Details
Domain Distribution.
The terminology in the dataset spans various AI-related domains. As shown in Figure 8, the six most frequent domains include Statistics and Probability (13.31%), Math (12.24%), Computer Science (11.74%), Natural Language Processing (11.50%), Data Science (9.98%), and Computer Vision (6.57%). The largest proportion of terms (34.65%) falls under the “Other” category, which represents interdisciplinary or less-defined concepts that do not fit neatly into any of the predefined categories. This distribution reflects the diversity and multidisciplinary nature of GIST.
Semantic Clustering.
We used Uniform Manifold Approximation and Projection (UMAP) to visualize the semantic relationships between terms across domains. Figure 9 shows a low-dimensional embedding of terms, where each point represents a term, color-coded by its domain. The visualization reveals distinct clusters corresponding to each domain, indicating strong intra-domain coherence. Overlaps between clusters (e.g., Data Science and Natural Language Processing) highlight the interconnected nature of these fields, where concepts are often shared or applied across domains.
Coverage Analysis through Rarefaction.
To assess the comprehensiveness of our terminology collection, we conducted a rarefaction analysis, which is visualized in Figure 10. This analysis reveals the relationship between the sample size of papers and the coverage of AI terminology. The curve demonstrates a characteristic asymptotic behavior, starting with a steep increase in coverage ratio from approximately 0.3 at 10% of papers to 0.6 at 30% of papers. As the subset size increases, the marginal gain in coverage gradually diminishes, reaching a coverage ratio of approximately 0.9 at 100% of the papers. The error bars, representing standard deviation across 50 random samples for each subset size, notably decrease as the sample size increases, indicating more stable coverage at larger sample sizes. The asymptotic nature of the curve approaching 0.9 coverage suggests that our current collection has achieved a robust representation of commonly used AI terminology, with additional papers likely to introduce increasingly specialized or niche terms at a decreasing rate. The relatively small error bars at larger sample sizes (80-100%) indicate high consistency in terminology coverage across different subsets of the literature, supporting the reliability of our collection methodology.
Figure 8: Terminology distribution of the top-6 AI domains in GIST.
Figure 9: UMap visualization of terms in GIST by domain.
Figure 10: Rarefaction curve showing the mean coverage ratio of AI terminology with respect to the subset size of papers. Each subset size is sampled 50 times, and the error bars represent the standard deviation of the coverage ratios across these samples.
A.6 Terminology Examples
To identify and show the most frequently used terms in the original set of awarded papers, we extracted and ranked the top 150 terminologies based on their occurrence frequency. Table 9 provides a comprehensive list of these terms, categorized by their rank and grouped for clarity.
Rank Terminology Rank1 Terminology (Continued)1 Rank2 Terminology (Continued)2
1 Algorithm 51 Latent space 101 Recall
2 Model 52 Node 102 Q-learning
3 Classifier 53 Transfer learning 103 Lasso
4 Transformer 54 Stochastic gradient descent 104 Transition matrix
5 Machine learning 55 Feature vector 105 Linear regression
6 Policy 56 Gibbs sampling 106 Meta-learning
7 Learning rate 57 Baseline 107 Segmentation
8 Neural network 58 Generative model 108 Fourier transform
9 Language model 59 Ontology 109 Epoch
10 Loss function 60 Attention 110 Learning algorithm
11 Reinforcement learning 61 Training set 111 Topic model
12 Encoder 62 Data mining 112 Time complexity
13 Deep learning 63 Manifold 113 Feature selection
14 Decoder 64 Discriminator 114 Knowledge distillation
15 Gradient descent 65 F1 score 115 Word embedding
16 Beam search 66 Dynamic programming 116 Euclidean distance
17 Machine translation 67 Adam optimizer 117 Covariance
18 Computer vision 68 Eigenvalue 118 Hyper-parameter
19 Dataset 69 Vector 119 Test set
20 Graph 70 State-of-the-art 120 Attention mechanism
21 Markov chain 71 Regularization 121 Oracle
22 Kernel 72 Backpropagation 122 Question answering
23 Marginal likelihood 73 Greedy algorithm 123 Point cloud
24 Objective function 74 Optical flow 124 Local minima
25 Gradient 75 Mutual information 125 N-gram
26 Reward function 76 Weight decay 126 Semi-supervised learning
27 Entropy 77 Posterior distribution 127 Batch normalization
28 Tensor 78 Bounding box 128 Homomorphism
29 Active learning 79 Disentanglement 129 Markov
30 Natural language processing 80 Convolution 130 Mini-batch
31 Perplexity 81 Semantic segmentation 131 Subgraph
32 Posterior 82 Logit 132 Bias
33 Logistic regression 83 Loss 133 Arg min
34 Covariance matrix 84 Multi-task learning 134 State space
35 Self-attention 85 Matrix 135 Dimensionality
36 Data augmentation 86 Binary classification 136 Random variable
37 Object detection 87 In-context learning 137 Gaussian distribution
38 Inference 88 Validation set 138 Optimizer
39 Cosine similarity 89 Cost function 139 Weight vector
40 Sample complexity 90 Corpus 140 Named entity recognition
41 Value function 91 Estimator 141 Kernel matrix
42 Probability distribution 92 Lemma 142 Discount factor
43 Generator 93 Parser 143 Hidden layer
44 Adam 94 Feature space 144 Domain adaptation
45 Supervised learning 95 Sentiment analysis 145 Frobenius norm
46 Dropout 96 Token 146 Positional encoding
47 Classification 97 Unsupervised learning 147 Seq2seq
48 Ground truth 98 State 148 Cross validation
49 Arg max 99 Infoset 149 Gaussian process
50 K-means 100 Precision 150 Coreference resolution
Table 9: Top 150 terms with the highest frequency in the original set of awarded papers.
Appendix B Additional Experiment Details
B.1 Terminology Refinement via Prompting Details
Figure 11: Prompts for refining a machine translation model’s initial output using relevant term translations from GIST as input.
In this subsection, we describe the process of refining machine translation outputs by leveraging prompts tailored for terminology consistency. The objective is to ensure that specific domain-related terms are translated accurately, adhering to predefined mappings provided in the dataset.
As illustrated in Figure 11, the prompt incorporates a term dictionary, the source language src_lang text, and the initial machine-translated tgt_lang output. The model is instructed to revise the translation by applying the specified target terms for corresponding source terms while maintaining the rest of the content unchanged. This approach ensures consistency and accuracy in translation outputs, particularly in specialized fields, by integrating domain knowledge directly into the model’s refinement process.
Appendix C Additional Experiment Results
We present comprehensive evaluation results using BLEU (Papineni et al., 2002), ChrF (Popović, 2015), ChrF++ (Popović, 2017), and TER (Snover et al., 2006) for translation into the five target languages across both evaluation sets.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 22.38 23.75 23.02
Bleu score std 24.43 25.89 24.92
Bleu 1gram precision mean 22.48 23.85 23.13
Bleu 1gram precision std 24.47 25.92 24.96
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 70.18 70.31 70.33
Bleu ref len std 35.62 35.60 35.84
Bleu pred len mean 70.08 70.24 70.17
Bleu pred len std 35.40 35.41 35.62
Bleu bp mean 0.99 0.99 0.99
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00
Bleu ratio std 0.02 0.02 0.02
Chrf score mean 22.87 24.24 23.51
Chrf score std 24.56 26.02 25.04
Ter score mean 91.12 89.53 90.37
Ter score std 29.39 31.08 29.94
Chrfpp score mean 22.87 24.24 23.51
Chrfpp score std 24.56 26.02 25.04
Table 10: BLEU, ChrF, ChrF++, and TER results for Arabic on the 60-60 evaluation set using nllb-200-3.3B.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 23.13 24.28 23.10
Bleu score std 24.72 25.84 24.57
Bleu 1gram precision mean 23.19 24.35 23.18
Bleu 1gram precision std 24.72 25.85 24.59
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 69.56 69.79 69.84
Bleu ref len std 35.32 35.54 35.56
Bleu pred len mean 69.64 69.89 69.83
Bleu pred len std 35.11 35.31 35.30
Bleu bp mean 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00
Bleu ratio std 0.02 0.02 0.02
Chrf score mean 23.67 24.83 23.64
Chrf score std 24.92 26.04 24.78
Ter score mean 90.27 88.94 90.29
Ter score std 29.63 31.00 29.45
Chrfpp score mean 23.67 24.83 23.64
Chrfpp score std 24.92 26.04 24.78
Table 11: BLEU, ChrF, ChrF++, and TER results for Arabic on the 60-60 evaluation set using hf-seamless-m4t-large.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 23.58 24.64 23.57
Bleu score std 26.63 27.85 26.40
Bleu 1gram precision mean 23.62 24.69 23.64
Bleu 1gram precision std 26.64 27.86 26.43
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 70.40 70.59 70.67
Bleu ref len std 35.88 35.95 36.14
Bleu pred len mean 70.68 70.85 70.88
Bleu pred len std 35.92 36.00 36.14
Bleu bp mean 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.01 1.00 1.00
Bleu ratio std 0.02 0.02 0.02
Chrf score mean 24.06 25.13 24.06
Chrf score std 26.76 27.98 26.55
Ter score mean 89.66 88.49 89.54
Ter score std 32.09 33.43 31.74
Chrfpp score mean 24.06 25.13 24.06
Chrfpp score std 26.76 27.98 26.55
Table 12: BLEU, ChrF, ChrF++, and TER results for Arabic on the 60-60 evaluation set using gpt-4o-mini.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 19.98 20.52 19.77
Bleu score std 22.35 23.49 22.14
Bleu 1gram precision mean 20.03 20.57 19.82
Bleu 1gram precision std 22.37 23.50 22.15
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 69.72 69.93 69.91
Bleu ref len std 35.74 36.02 35.93
Bleu pred len mean 69.99 70.20 70.14
Bleu pred len std 35.69 35.97 35.85
Bleu bp mean 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.01 1.01 1.01
Bleu ratio std 0.02 0.02 0.02
Chrf score mean 20.50 21.03 20.28
Chrf score std 22.54 23.65 22.34
Ter score mean 93.97 93.40 94.20
Ter score std 26.96 28.24 26.69
Chrfpp score mean 20.50 21.03 20.28
Chrfpp score std 22.54 23.65 22.34
Table 13: BLEU, ChrF, ChrF++, and TER results for Arabic on the 60-60 evaluation set using aya-23-8B.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 20.11 21.34 20.29
Bleu score std 23.09 24.68 23.43
Bleu 1gram precision mean 20.20 21.43 20.38
Bleu 1gram precision std 23.14 24.73 23.49
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 71.54 71.47 71.70
Bleu ref len std 36.10 36.29 36.28
Bleu pred len mean 71.79 71.75 71.93
Bleu pred len std 36.24 36.43 36.39
Bleu bp mean 0.99 0.99 0.99
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00
Bleu ratio std 0.02 0.02 0.02
Chrf score mean 20.62 21.85 20.80
Chrf score std 23.26 24.83 23.60
Ter score mean 93.61 92.27 93.37
Ter score std 27.75 29.58 28.12
Chrfpp score mean 20.62 21.85 20.80
Chrfpp score std 23.26 24.83 23.60
Table 14: BLEU, ChrF, ChrF++, and TER results for Arabic on the 60-60 evaluation set using aya-expanse-8B.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 17.29 19.21 18.31
Bleu score std 22.88 24.37 24.26
Bleu 1gram precision mean 17.37 19.30 18.40
Bleu 1gram precision std 22.94 24.43 24.32
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 24.74 26.52 24.75
Bleu ref len std 12.69 13.49 12.72
Bleu pred len mean 24.59 26.35 24.60
Bleu pred len std 12.46 13.22 12.49
Bleu bp mean 0.99 0.99 0.99
Bleu bp std 0.02 0.02 0.02
Bleu ratio mean 1.00 1.00 1.00
Bleu ratio std 0.03 0.03 0.03
Chrf score mean 17.45 19.36 18.47
Chrf score std 23.08 24.54 24.43
Ter score mean 83.26 81.33 82.23
Ter score std 23.01 24.53 24.40
Chrfpp score mean 17.45 19.36 18.47
Chrfpp score std 23.08 24.54 24.43
Table 15: BLEU, ChrF, ChrF++, and TER results for Chinese on the 60-60 evaluation set using nllb-200-3.3B.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 26.26 27.23 27.06
Bleu score std 24.84 26.18 25.59
Bleu 1gram precision mean 26.51 27.50 27.32
Bleu 1gram precision std 24.89 26.23 25.65
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 28.70 29.35 28.76
Bleu ref len std 14.62 14.70 14.65
Bleu pred len mean 28.24 28.86 28.30
Bleu pred len std 13.91 13.97 13.96
Bleu bp mean 0.99 0.99 0.99
Bleu bp std 0.03 0.03 0.03
Bleu ratio mean 0.99 0.99 0.99
Bleu ratio std 0.04 0.04 0.04
Chrf score mean 26.43 27.41 27.23
Chrf score std 24.92 26.25 25.66
Ter score mean 74.32 73.34 73.51
Ter score std 25.21 26.56 25.97
Chrfpp score mean 26.43 27.41 27.23
Chrfpp score std 24.92 26.25 25.66
Table 16: BLEU, ChrF, ChrF++, and TER results for Chinese on the 60-60 evaluation set using hf-seamless-m4t-large.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 32.64 34.24 33.30
Bleu score std 29.93 31.10 30.52
Bleu 1gram precision mean 32.69 34.30 33.34
Bleu 1gram precision std 29.94 31.12 30.53
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 29.97 30.02 29.94
Bleu ref len std 14.74 14.79 14.77
Bleu pred len mean 30.02 30.06 30.00
Bleu pred len std 14.73 14.74 14.77
Bleu bp mean 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00
Bleu ratio std 0.02 0.02 0.02
Chrf score mean 32.76 34.36 33.42
Chrf score std 29.94 31.11 30.53
Ter score mean 67.79 66.19 67.13
Ter score std 30.20 31.39 30.79
Chrfpp score mean 32.76 34.36 33.42
Chrfpp score std 29.94 31.11 30.53
Table 17: BLEU, ChrF, ChrF++, and TER results for Chinese on the 60-60 evaluation set using gpt-4o-mini.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 26.08 26.55 26.47
Bleu score std 25.69 26.49 25.63
Bleu 1gram precision mean 26.17 26.64 26.56
Bleu 1gram precision std 25.73 26.53 25.67
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 29.33 29.36 29.31
Bleu ref len std 14.47 14.48 14.52
Bleu pred len mean 29.25 29.28 29.23
Bleu pred len std 14.24 14.24 14.29
Bleu bp mean 1.00 1.00 1.00
Bleu bp std 0.02 0.02 0.02
Bleu ratio mean 1.00 1.00 1.00
Bleu ratio std 0.03 0.03 0.03
Chrf score mean 26.20 26.66 26.59
Chrf score std 25.72 26.52 25.66
Ter score mean 74.40 73.94 74.02
Ter score std 25.99 26.77 25.91
Chrfpp score mean 26.20 26.66 26.59
Chrfpp score std 25.72 26.52 25.66
Table 18: BLEU, ChrF, ChrF++, and TER results for Chinese on the 60-60 evaluation set using aya-23-8B.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 27.31 28.63 27.55
Bleu score std 27.56 28.76 27.73
Bleu 1gram precision mean 27.47 28.75 27.71
Bleu 1gram precision std 27.61 28.80 27.78
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 30.61 30.47 30.57
Bleu ref len std 14.86 14.93 14.87
Bleu pred len mean 30.54 30.46 30.50
Bleu pred len std 14.81 14.83 14.82
Bleu bp mean 0.99 0.99 0.99
Bleu bp std 0.04 0.02 0.04
Bleu ratio mean 1.00 1.00 1.00
Bleu ratio std 0.04 0.03 0.04
Chrf score mean 27.52 28.83 27.76
Chrf score std 27.69 28.88 27.86
Ter score mean 73.47 72.14 73.22
Ter score std 27.92 29.14 28.10
Chrfpp score mean 27.52 28.83 27.76
Chrfpp score std 27.69 28.88 27.86
Table 19: BLEU, ChrF, ChrF++, and TER results for Chinese on the 60-60 evaluation set using aya-expanse-8B.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 34.93 37.78 35.14
Bleu score std 29.70 31.33 29.69
Bleu 1gram precision mean 35.10 37.91 35.31
Bleu 1gram precision std 29.76 31.37 29.75
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 100.59 101.31 101.12
Bleu ref len std 48.91 49.80 49.06
Bleu pred len mean 100.42 101.37 100.88
Bleu pred len std 48.75 49.74 48.84
Bleu bp mean 0.99 1.00 0.99
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00
Bleu ratio std 0.02 0.02 0.02
Chrf score mean 35.55 38.40 35.76
Chrf score std 29.85 31.44 29.83
Ter score mean 75.21 71.97 74.95
Ter score std 34.89 36.82 34.81
Chrfpp score mean 35.55 38.40 35.76
Chrfpp score std 29.85 31.44 29.83
Table 20: BLEU, ChrF, ChrF++, and TER results for French on the 60-60 evaluation set using nllb-200-3.3B.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 40.04 42.12 39.47
Bleu score std 31.35 32.61 31.61
Bleu 1gram precision mean 40.17 42.22 39.60
Bleu 1gram precision std 31.38 32.63 31.64
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 101.49 101.82 101.87
Bleu ref len std 50.11 50.22 50.16
Bleu pred len mean 101.56 101.99 101.88
Bleu pred len std 49.98 50.12 49.98
Bleu bp mean 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00
Bleu ratio std 0.02 0.02 0.02
Chrf score mean 40.68 42.76 40.10
Chrf score std 31.39 32.64 31.65
Ter score mean 69.19 66.81 69.85
Ter score std 36.78 38.18 37.00
Chrfpp score mean 40.68 42.76 40.10
Chrfpp score std 31.39 32.64 31.65
Table 21: BLEU, ChrF, ChrF++, and TER results for French on the 60-60 evaluation set using hf-seamless-m4t-large.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 40.80 43.88 41.30
Bleu score std 32.29 33.65 32.31
Bleu 1gram precision mean 40.90 43.98 41.41
Bleu 1gram precision std 32.30 33.67 32.32
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 102.43 102.86 102.68
Bleu ref len std 49.81 50.12 49.94
Bleu pred len mean 102.54 102.98 102.72
Bleu pred len std 49.77 50.05 49.84
Bleu bp mean 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00
Bleu ratio std 0.02 0.02 0.02
Chrf score mean 41.40 44.46 41.89
Chrf score std 32.27 33.61 32.28
Ter score mean 68.32 64.76 67.77
Ter score std 37.85 39.39 37.85
Chrfpp score mean 41.40 44.46 41.89
Chrfpp score std 32.27 33.61 32.28
Table 22: BLEU, ChrF, ChrF++, and TER results for French on the 60-60 evaluation set using gpt-4o-mini.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 33.85 36.13 33.74
Bleu score std 28.54 30.46 28.57
Bleu 1gram precision mean 33.98 36.23 33.87
Bleu 1gram precision std 28.57 30.48 28.60
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 100.78 101.15 101.19
Bleu ref len std 49.88 49.82 49.92
Bleu pred len mean 100.70 101.22 101.07
Bleu pred len std 49.64 49.64 49.61
Bleu bp mean 0.99 1.00 0.99
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00
Bleu ratio std 0.02 0.02 0.02
Chrf score mean 34.48 36.74 34.37
Chrf score std 28.63 30.53 28.66
Ter score mean 76.35 73.71 76.53
Ter score std 33.54 35.72 33.57
Chrfpp score mean 34.48 36.74 34.37
Chrfpp score std 28.63 30.53 28.66
Table 23: BLEU, ChrF, ChrF++, and TER results for French on the 60-60 evaluation set using aya-23-8B.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 33.05 35.50 33.24
Bleu score std 28.46 30.64 28.72
Bleu 1gram precision mean 33.18 35.62 33.38
Bleu 1gram precision std 28.50 30.67 28.75
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 102.37 102.52 102.66
Bleu ref len std 49.92 50.07 49.96
Bleu pred len mean 102.37 102.62 102.62
Bleu pred len std 49.76 49.92 49.77
Bleu bp mean 0.99 1.00 0.99
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00
Bleu ratio std 0.02 0.02 0.02
Chrf score mean 33.68 36.12 33.86
Chrf score std 28.56 30.71 28.81
Ter score mean 77.43 74.65 77.20
Ter score std 33.50 36.00 33.75
Chrfpp score mean 33.68 36.12 33.86
Chrfpp score std 28.56 30.71 28.81
Table 24: BLEU, ChrF, ChrF++, and TER results for French on the 60-60 evaluation set using aya-expanse-8B.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 6.19 8.61 6.72
Bleu score std 12.90 16.15 13.59
Bleu 1gram precision mean 6.26 8.67 6.79
Bleu 1gram precision std 12.94 16.17 13.64
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 36.22 38.30 36.51
Bleu ref len std 20.17 20.98 20.20
Bleu pred len mean 35.45 37.76 35.74
Bleu pred len std 19.73 20.62 19.75
Bleu bp mean 0.98 0.98 0.98
Bleu bp std 0.03 0.03 0.03
Bleu ratio mean 0.98 0.99 0.98
Bleu ratio std 0.04 0.04 0.04
Chrf score mean 6.22 8.64 6.75
Chrf score std 12.93 16.17 13.62
Ter score mean 94.32 91.91 93.79
Ter score std 13.12 16.39 13.81
Chrfpp score mean 6.22 8.64 6.75
Chrfpp score std 12.93 16.17 13.62
Table 25: BLEU, ChrF, ChrF++, and TER results for Japanese on the 60-60 evaluation set using nllb-200-3.3B.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 14.56 15.30 14.61
Bleu score std 20.15 20.98 19.70
Bleu 1gram precision mean 14.71 15.43 14.76
Bleu 1gram precision std 20.30 21.10 19.86
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 40.61 40.93 40.74
Bleu ref len std 19.85 20.05 19.91
Bleu pred len mean 40.24 40.66 40.37
Bleu pred len std 19.70 19.97 19.76
Bleu bp mean 0.99 0.99 0.99
Bleu bp std 0.02 0.02 0.02
Bleu ratio mean 0.99 0.99 0.99
Bleu ratio std 0.03 0.03 0.03
Chrf score mean 14.65 15.39 14.70
Chrf score std 20.22 21.05 19.77
Ter score mean 86.02 85.28 85.97
Ter score std 20.37 21.21 19.92
Chrfpp score mean 14.65 15.39 14.70
Chrfpp score std 20.22 21.05 19.77
Table 26: BLEU, ChrF, ChrF++, and TER results for Japanese on the 60-60 evaluation set using hf-seamless-m4t-large.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 21.46 22.11 21.66
Bleu score std 23.81 24.77 23.93
Bleu 1gram precision mean 21.47 22.11 21.66
Bleu 1gram precision std 23.81 24.77 23.93
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 43.57 43.54 43.59
Bleu ref len std 21.17 21.10 21.08
Bleu pred len mean 43.88 43.85 43.90
Bleu pred len std 21.21 21.14 21.11
Bleu bp mean 1.00 1.00 1.00
Bleu bp std 0.00 0.00 0.00
Bleu ratio mean 1.01 1.01 1.01
Bleu ratio std 0.02 0.02 0.02
Chrf score mean 21.59 22.24 21.79
Chrf score std 23.89 24.85 24.02
Ter score mean 79.33 78.68 79.13
Ter score std 24.23 25.20 24.36
Chrfpp score mean 21.59 22.24 21.79
Chrfpp score std 23.89 24.85 24.02
Table 27: BLEU, ChrF, ChrF++, and TER results for Japanese on the 60-60 evaluation set using gpt-4o-mini.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 15.06 15.94 15.43
Bleu score std 18.77 19.65 18.65
Bleu 1gram precision mean 15.07 15.95 15.44
Bleu 1gram precision std 18.77 19.66 18.65
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 42.55 42.76 42.64
Bleu ref len std 21.06 21.06 21.01
Bleu pred len mean 42.81 43.01 42.90
Bleu pred len std 21.03 21.03 20.98
Bleu bp mean 1.00 1.00 1.00
Bleu bp std 0.00 0.01 0.00
Bleu ratio mean 1.01 1.01 1.01
Bleu ratio std 0.02 0.02 0.02
Chrf score mean 15.22 16.10 15.59
Chrf score std 19.06 19.92 18.94
Ter score mean 85.73 84.86 85.37
Ter score std 19.01 19.92 18.90
Chrfpp score mean 15.22 16.10 15.59
Chrfpp score std 19.06 19.92 18.94
Table 28: BLEU, ChrF, ChrF++, and TER results for Japanese on the 60-60 evaluation set using aya-23-8B.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 14.59 15.21 14.92
Bleu score std 18.20 18.91 18.42
Bleu 1gram precision mean 14.60 15.21 14.92
Bleu 1gram precision std 18.20 18.91 18.42
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 43.77 43.84 43.86
Bleu ref len std 21.07 21.10 21.02
Bleu pred len mean 44.09 44.16 44.18
Bleu pred len std 21.07 21.09 21.02
Bleu bp mean 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.01 1.01 1.01
Bleu ratio std 0.02 0.02 0.02
Chrf score mean 14.76 15.37 15.08
Chrf score std 18.48 19.18 18.70
Ter score mean 86.42 85.81 86.10
Ter score std 18.55 19.28 18.78
Chrfpp score mean 14.76 15.37 15.08
Chrfpp score std 18.48 19.18 18.70
Table 29: BLEU, ChrF, ChrF++, and TER results for Japanese on the 60-60 evaluation set using aya-expanse-8B.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 17.30 18.84 18.37
Bleu score std 21.32 22.75 22.14
Bleu 1gram precision mean 17.38 18.92 18.46
Bleu 1gram precision std 21.36 22.79 22.19
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 90.80 91.27 90.87
Bleu ref len std 45.14 45.83 45.23
Bleu pred len mean 90.83 91.38 90.85
Bleu pred len std 45.03 45.77 45.10
Bleu bp mean 0.99 0.99 0.99
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00
Bleu ratio std 0.03 0.03 0.03
Chrf score mean 17.69 19.24 18.77
Chrf score std 21.48 22.89 22.31
Ter score mean 93.87 92.19 92.70
Ter score std 24.61 26.24 25.62
Chrfpp score mean 17.69 19.24 18.77
Chrfpp score std 21.48 22.89 22.31
Table 30: BLEU, ChrF, ChrF++, and TER results for Russian on the 60-60 evaluation set using nllb-200-3.3B.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 17.18 18.88 18.34
Bleu score std 19.81 21.88 21.12
Bleu 1gram precision mean 17.24 18.95 18.41
Bleu 1gram precision std 19.84 21.91 21.15
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 91.72 91.87 91.81
Bleu ref len std 46.73 47.06 46.83
Bleu pred len mean 91.96 92.12 91.98
Bleu pred len std 46.59 46.92 46.68
Bleu bp mean 1.00 1.00 0.99
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.01 1.01 1.01
Bleu ratio std 0.03 0.03 0.03
Chrf score mean 17.61 19.33 18.79
Chrf score std 20.07 22.12 21.37
Ter score mean 94.01 92.13 92.71
Ter score std 22.88 25.28 24.39
Chrfpp score mean 17.61 19.33 18.79
Chrfpp score std 20.07 22.12 21.37
Table 31: BLEU, ChrF, ChrF++, and TER results for Russian on the 60-60 evaluation set using hf-seamless-m4t-large.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 17.25 18.33 17.12
Bleu score std 20.92 21.76 20.17
Bleu 1gram precision mean 17.32 18.39 17.19
Bleu 1gram precision std 20.96 21.80 20.23
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 92.78 92.61 92.73
Bleu ref len std 46.68 46.69 46.77
Bleu pred len mean 93.00 92.83 92.89
Bleu pred len std 46.56 46.58 46.62
Bleu bp mean 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.01 1.00 1.00
Bleu ratio std 0.02 0.02 0.02
Chrf score mean 17.65 18.73 17.52
Chrf score std 21.10 21.95 20.38
Ter score mean 93.90 92.73 94.05
Ter score std 24.29 25.30 23.46
Chrfpp score mean 17.65 18.73 17.52
Chrfpp score std 21.10 21.95 20.38
Table 32: BLEU, ChrF, ChrF++, and TER results for Russian on the 60-60 evaluation set using gpt-4o-mini.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 15.77 16.83 16.14
Bleu score std 19.10 20.61 19.55
Bleu 1gram precision mean 15.84 16.90 16.22
Bleu 1gram precision std 19.13 20.64 19.59
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 91.71 91.64 91.62
Bleu ref len std 46.58 46.43 46.64
Bleu pred len mean 91.83 91.76 91.69
Bleu pred len std 46.37 46.23 46.42
Bleu bp mean 0.99 0.99 0.99
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00
Bleu ratio std 0.03 0.03 0.03
Chrf score mean 16.17 17.23 16.54
Chrf score std 19.31 20.81 19.76
Ter score mean 95.60 94.42 95.15
Ter score std 22.20 23.92 22.73
Chrfpp score mean 16.17 17.23 16.54
Chrfpp score std 19.31 20.81 19.76
Table 33: BLEU, ChrF, ChrF++, and TER results for Russian on the 60-60 evaluation set using aya-23-8B.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 16.59 18.18 16.54
Bleu score std 20.30 22.44 20.48
Bleu 1gram precision mean 16.67 18.26 16.62
Bleu 1gram precision std 20.32 22.47 20.50
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 94.17 93.52 94.01
Bleu ref len std 46.56 46.42 46.60
Bleu pred len mean 94.25 93.63 94.07
Bleu pred len std 46.48 46.33 46.54
Bleu bp mean 0.99 0.99 0.99
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00
Bleu ratio std 0.03 0.02 0.02
Chrf score mean 16.99 18.57 16.92
Chrf score std 20.48 22.59 20.64
Ter score mean 94.56 92.82 94.64
Ter score std 23.46 25.92 23.68
Chrfpp score mean 16.99 18.57 16.92
Chrfpp score std 20.48 22.59 20.64
Table 34: BLEU, ChrF, ChrF++, and TER results for Russian on the 60-60 evaluation set using aya-expanse-8B.
Direct Translation Prompting Refinement Word Alignment
nllb-200-3.3B
Arabic 0.84 0.84 0.83
Chinese 0.78 0.81 0.79
French 0.83 0.84 0.83
Japanese 0.78 0.82 0.78
Russian 0.81 0.82 0.80
hf-seamless-m4t-large
Arabic 0.84 0.85 0.84
Chinese 0.83 0.85 0.84
French 0.84 0.85 0.84
Japanese 0.85 0.86 0.85
Russian 0.82 0.83 0.80
gpt-4o-mini
Arabic 0.86 0.86 0.85
Chinese 0.87 0.88 0.88
French 0.85 0.85 0.85
Japanese 0.90 0.90 0.90
Russian 0.84 0.84 0.81
aya-23-8B
Arabic 0.84 0.85 0.84
Chinese 0.85 0.86 0.85
French 0.82 0.83 0.82
Japanese 0.88 0.88 0.88
Russian 0.82 0.82 0.80
aya-expanse-8B
Arabic 0.82 0.83 0.81
Chinese 0.83 0.85 0.84
French 0.82 0.83 0.82
Japanese 0.89 0.89 0.89
Russian 0.82 0.83 0.80
Table 35: COMET results on the 60-60 evaluation set.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 15.42 15.12 14.66
Bleu score std 15.76 16.02 14.65
Bleu 1gram precision mean 15.54 15.21 14.76
Bleu 1gram precision std 15.83 16.06 14.68
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 169.01 177.84 176.63
Bleu ref len std 65.20 64.39 63.91
Bleu pred len mean 168.72 177.87 176.63
Bleu pred len std 64.71 64.41 64.06
Bleu bp mean 0.99 0.99 0.99
Bleu bp std 0.03 0.02 0.02
Bleu ratio mean 1.00 1.00 1.00
Bleu ratio std 0.04 0.03 0.03
Chrf score mean 15.95 15.64 15.18
Chrf score std 16.00 16.32 14.93
Ter score mean 101.29 101.47 102.00
Ter score std 20.20 20.06 18.57
Chrfpp score mean 15.95 15.64 15.18
Chrfpp score std 16.00 16.32 14.93
Table 36: BLEU, ChrF, ChrF++, and TER results for Arabic on the AI papers and model cards evaluation set using nllb-200-3.3B.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 15.38 16.48 15.83
Bleu score std 16.32 17.52 16.01
Bleu 1gram precision mean 15.50 16.58 15.93
Bleu 1gram precision std 16.40 17.58 16.06
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 173.42 180.00 179.57
Bleu ref len std 66.22 64.01 63.90
Bleu pred len mean 173.41 179.91 179.47
Bleu pred len std 66.22 64.04 63.99
Bleu bp mean 0.99 0.99 0.99
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00
Bleu ratio std 0.03 0.02 0.02
Chrf score mean 15.90 17.00 16.36
Chrf score std 16.57 17.76 16.25
Ter score mean 101.34 99.86 100.57
Ter score std 21.12 21.79 19.97
Chrfpp score mean 15.90 17.00 16.36
Chrfpp score std 16.57 17.76 16.25
Table 37: BLEU, ChrF, ChrF++, and TER results for Arabic on the AI papers and model cards evaluation set using hf-seamless-m4t-large.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 14.37 14.90 13.95
Bleu score std 16.32 15.58 14.30
Bleu 1gram precision mean 14.42 14.95 14.00
Bleu 1gram precision std 16.32 15.59 14.32
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 174.18 181.13 181.07
Bleu ref len std 67.36 65.07 65.15
Bleu pred len mean 174.26 181.55 181.37
Bleu pred len std 67.26 65.19 65.31
Bleu bp mean 0.99 1.00 1.00
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00
Bleu ratio std 0.02 0.01 0.02
Chrf score mean 14.82 15.37 14.40
Chrf score std 16.47 15.76 14.49
Ter score mean 102.33 101.29 102.58
Ter score std 20.70 19.21 17.92
Chrfpp score mean 14.82 15.37 14.40
Chrfpp score std 16.47 15.76 14.49
Table 38: BLEU, ChrF, ChrF++, and TER results for Arabic on the AI papers and model cards evaluation set using gpt-4o-mini.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 14.28 15.08 14.55
Bleu score std 14.76 16.23 15.04
Bleu 1gram precision mean 14.34 15.15 14.63
Bleu 1gram precision std 14.82 16.29 15.11
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 177.82 179.16 178.20
Bleu ref len std 65.99 65.72 66.07
Bleu pred len mean 178.21 179.61 178.54
Bleu pred len std 66.06 65.78 66.13
Bleu bp mean 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00
Bleu ratio std 0.02 0.02 0.02
Chrf score mean 14.77 15.59 15.05
Chrf score std 15.02 16.49 15.29
Ter score mean 102.45 101.53 102.09
Ter score std 18.88 20.35 18.95
Chrfpp score mean 14.77 15.59 15.05
Chrfpp score std 15.02 16.49 15.29
Table 39: BLEU, ChrF, ChrF++, and TER results for Arabic on the AI papers and model cards evaluation set using aya-23-8B.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 11.47 11.83 11.57
Bleu score std 10.75 11.92 11.66
Bleu 1gram precision mean 11.50 11.86 11.61
Bleu 1gram precision std 10.78 11.94 11.69
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 183.01 183.18 183.27
Bleu ref len std 66.11 66.21 66.10
Bleu pred len mean 185.58 185.77 185.77
Bleu pred len std 66.51 66.57 66.44
Bleu bp mean 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.02 1.02 1.02
Bleu ratio std 0.05 0.05 0.05
Chrf score mean 11.95 12.32 12.05
Chrf score std 11.00 12.15 11.88
Ter score mean 106.35 106.05 106.22
Ter score std 15.12 16.63 15.96
Chrfpp score mean 11.95 12.32 12.05
Chrfpp score std 11.00 12.15 11.88
Table 40: BLEU, ChrF, ChrF++, and TER results for Arabic on the AI papers and model cards evaluation set using aya-expanse-8B.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 10.24 12.43 12.31
Bleu score std 11.68 14.39 13.71
Bleu 1gram precision mean 10.29 12.51 12.38
Bleu 1gram precision std 11.72 14.45 13.76
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 75.92 80.18 79.39
Bleu ref len std 31.48 32.63 32.60
Bleu pred len mean 79.72 84.07 83.38
Bleu pred len std 34.40 35.19 35.41
Bleu bp mean 0.98 0.99 0.99
Bleu bp std 0.11 0.08 0.09
Bleu ratio mean 1.04 1.05 1.05
Bleu ratio std 0.14 0.12 0.13
Chrf score mean 10.82 13.08 12.98
Chrf score std 12.05 14.82 14.16
Ter score mean 101.05 98.60 98.46
Ter score std 19.34 21.01 19.97
Chrfpp score mean 10.82 13.08 12.98
Chrfpp score std 12.05 14.82 14.16
Table 41: BLEU, ChrF, ChrF++, and TER results for Chinese on the AI papers and model cards evaluation set using nllb-200-3.3B.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 13.67 14.77 14.40
Bleu score std 15.29 15.79 16.04
Bleu 1gram precision mean 13.75 14.86 14.48
Bleu 1gram precision std 15.31 15.84 16.08
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 75.67 81.10 78.32
Bleu ref len std 34.88 34.04 33.85
Bleu pred len mean 78.80 84.24 81.61
Bleu pred len std 36.63 35.88 35.88
Bleu bp mean 0.99 0.99 0.99
Bleu bp std 0.05 0.04 0.04
Bleu ratio mean 1.04 1.04 1.04
Bleu ratio std 0.08 0.07 0.07
Chrf score mean 14.34 15.42 15.06
Chrf score std 15.65 16.10 16.38
Ter score mean 97.26 95.50 96.19
Ter score std 22.32 21.46 22.40
Chrfpp score mean 14.34 15.42 15.06
Chrfpp score std 15.65 16.10 16.38
Table 42: BLEU, ChrF, ChrF++, and TER results for Chinese on the AI papers and model cards evaluation set using hf-seamless-m4t-large.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 17.21 18.44 18.60
Bleu score std 18.76 19.49 19.25
Bleu 1gram precision mean 17.35 18.54 18.70
Bleu 1gram precision std 18.95 19.62 19.38
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 83.21 85.58 85.50
Bleu ref len std 36.03 35.24 35.33
Bleu pred len mean 86.68 88.99 88.96
Bleu pred len std 37.81 37.00 37.10
Bleu bp mean 0.99 1.00 1.00
Bleu bp std 0.02 0.02 0.02
Bleu ratio mean 1.04 1.04 1.04
Bleu ratio std 0.06 0.06 0.06
Chrf score mean 17.93 19.16 19.32
Chrf score std 19.10 19.82 19.56
Ter score mean 93.01 90.78 90.55
Ter score std 24.89 24.17 24.00
Chrfpp score mean 17.93 19.16 19.32
Chrfpp score std 19.10 19.82 19.56
Table 43: BLEU, ChrF, ChrF++, and TER results for Chinese on the AI papers and model cards evaluation set using gpt-4o-mini.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 14.50 14.90 14.70
Bleu score std 15.36 16.07 15.52
Bleu 1gram precision mean 14.58 14.98 14.78
Bleu 1gram precision std 15.49 16.20 15.65
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 84.04 84.57 84.06
Bleu ref len std 35.54 35.28 35.58
Bleu pred len mean 87.66 88.19 87.69
Bleu pred len std 37.74 37.49 37.81
Bleu bp mean 1.00 1.00 1.00
Bleu bp std 0.02 0.02 0.02
Bleu ratio mean 1.04 1.04 1.04
Bleu ratio std 0.06 0.06 0.06
Chrf score mean 15.12 15.52 15.32
Chrf score std 15.65 16.35 15.81
Ter score mean 95.18 94.49 94.76
Ter score std 20.33 20.54 20.07
Chrfpp score mean 15.12 15.52 15.32
Chrfpp score std 15.65 16.35 15.81
Table 44: BLEU, ChrF, ChrF++, and TER results for Chinese on the AI papers and model cards evaluation set using aya-23-8B.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 12.04 12.98 12.21
Bleu score std 13.46 14.42 13.42
Bleu 1gram precision mean 12.08 13.03 12.24
Bleu 1gram precision std 13.51 14.47 13.46
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 86.46 86.23 86.48
Bleu ref len std 36.21 36.19 36.24
Bleu pred len mean 92.34 91.99 92.37
Bleu pred len std 39.90 39.68 39.94
Bleu bp mean 1.00 1.00 1.00
Bleu bp std 0.01 0.02 0.01
Bleu ratio mean 1.07 1.07 1.07
Bleu ratio std 0.08 0.09 0.08
Chrf score mean 12.76 13.74 12.94
Chrf score std 13.86 14.81 13.83
Ter score mean 99.34 98.29 99.13
Ter score std 20.49 21.52 20.54
Chrfpp score mean 12.76 13.74 12.94
Chrfpp score std 13.86 14.81 13.83
Table 45: BLEU, ChrF, ChrF++, and TER results for Chinese on the AI papers and model cards evaluation set using aya-expanse-8B.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 22.68 25.41 23.57
Bleu score std 20.79 23.06 21.92
Bleu 1gram precision mean 22.75 25.47 23.64
Bleu 1gram precision std 20.83 23.10 21.96
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 203.28 234.13 207.02
Bleu ref len std 84.53 89.52 84.94
Bleu pred len mean 204.82 235.82 208.57
Bleu pred len std 85.31 90.00 85.73
Bleu bp mean 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.01 1.01 1.01
Bleu ratio std 0.03 0.03 0.03
Chrf score mean 23.33 26.10 24.25
Chrf score std 21.07 23.29 22.15
Ter score mean 91.40 88.33 90.51
Ter score std 25.52 27.99 26.91
Chrfpp score mean 23.33 26.10 24.25
Chrfpp score std 21.07 23.29 22.15
Table 46: BLEU, ChrF, ChrF++, and TER results for French on the AI papers and model cards evaluation set using nllb-200-3.3B.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 24.34 29.55 25.83
Bleu score std 20.94 24.54 21.80
Bleu 1gram precision mean 24.42 29.60 25.90
Bleu 1gram precision std 20.98 24.57 21.84
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 237.17 246.51 243.30
Bleu ref len std 90.05 87.83 87.19
Bleu pred len mean 238.76 248.14 244.78
Bleu pred len std 91.05 88.54 87.84
Bleu bp mean 0.99 1.00 1.00
Bleu bp std 0.03 0.02 0.02
Bleu ratio mean 1.00 1.01 1.01
Bleu ratio std 0.03 0.02 0.03
Chrf score mean 25.05 30.30 26.55
Chrf score std 21.09 24.71 22.03
Ter score mean 89.75 83.28 87.62
Ter score std 26.10 29.48 26.32
Chrfpp score mean 25.05 30.30 26.55
Chrfpp score std 21.09 24.71 22.03
Table 47: BLEU, ChrF, ChrF++, and TER results for French on the AI papers and model cards evaluation set using hf-seamless-m4t-large.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 24.45 28.83 25.73
Bleu score std 22.85 26.47 23.16
Bleu 1gram precision mean 24.51 28.89 25.79
Bleu 1gram precision std 22.88 26.50 23.19
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 245.51 252.04 251.56
Bleu ref len std 96.46 91.08 91.27
Bleu pred len mean 245.92 252.59 251.95
Bleu pred len std 96.66 91.19 91.33
Bleu bp mean 0.99 1.00 1.00
Bleu bp std 0.03 0.02 0.02
Bleu ratio mean 1.00 1.00 1.00
Bleu ratio std 0.03 0.02 0.02
Chrf score mean 25.04 29.44 26.35
Chrf score std 22.96 26.59 23.33
Ter score mean 89.16 83.80 87.36
Ter score std 27.90 31.46 27.52
Chrfpp score mean 25.04 29.44 26.35
Chrfpp score std 22.96 26.59 23.33
Table 48: BLEU, ChrF, ChrF++, and TER results for French on the AI papers and model cards evaluation set using gpt-4o-mini.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 24.49 26.86 24.57
Bleu score std 21.92 23.53 22.03
Bleu 1gram precision mean 24.58 26.93 24.65
Bleu 1gram precision std 21.96 23.57 22.07
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 246.19 249.56 246.63
Bleu ref len std 93.04 90.93 93.09
Bleu pred len mean 246.42 250.05 246.80
Bleu pred len std 92.99 90.89 92.96
Bleu bp mean 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00
Bleu ratio std 0.02 0.02 0.02
Chrf score mean 25.11 27.48 25.18
Chrf score std 22.11 23.67 22.24
Ter score mean 89.07 86.14 88.79
Ter score std 26.62 28.31 26.49
Chrfpp score mean 25.11 27.48 25.18
Chrfpp score std 22.11 23.67 22.24
Table 49: BLEU, ChrF, ChrF++, and TER results for French on the AI papers and model cards evaluation set using aya-23-8B.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 18.84 20.56 17.99
Bleu score std 18.05 19.71 16.23
Bleu 1gram precision mean 18.89 20.60 18.04
Bleu 1gram precision std 18.10 19.75 16.27
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 252.31 252.42 252.40
Bleu ref len std 91.64 91.62 91.68
Bleu pred len mean 255.15 255.25 255.12
Bleu pred len std 91.11 91.18 91.09
Bleu bp mean 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.02 1.02 1.02
Bleu ratio std 0.05 0.05 0.05
Chrf score mean 19.49 21.23 18.65
Chrf score std 18.37 20.02 16.58
Ter score mean 96.12 94.10 97.10
Ter score std 22.84 24.64 20.72
Chrfpp score mean 19.49 21.23 18.65
Chrfpp score std 18.37 20.02 16.58
Table 50: BLEU, ChrF, ChrF++, and TER results for French on the AI papers and model cards evaluation set using aya-expanse-8B.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 8.24 9.33 9.13
Bleu score std 10.82 11.97 11.65
Bleu 1gram precision mean 8.33 9.41 9.22
Bleu 1gram precision std 10.95 12.10 11.79
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 108.20 113.09 112.40
Bleu ref len std 45.16 43.78 44.13
Bleu pred len mean 112.33 117.01 116.40
Bleu pred len std 47.46 46.19 46.72
Bleu bp mean 0.99 0.99 0.99
Bleu bp std 0.07 0.05 0.05
Bleu ratio mean 1.03 1.03 1.03
Bleu ratio std 0.10 0.08 0.09
Chrf score mean 8.61 9.72 9.53
Chrf score std 11.04 12.30 11.97
Ter score mean 101.41 100.10 100.54
Ter score std 16.93 16.77 16.69
Chrfpp score mean 8.61 9.72 9.53
Chrfpp score std 11.04 12.30 11.97
Table 51: BLEU, ChrF, ChrF++, and TER results for Japanese on the AI papers and model cards evaluation set using nllb-200-3.3B.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 9.42 9.98 9.84
Bleu score std 11.49 12.78 12.60
Bleu 1gram precision mean 9.48 10.03 9.89
Bleu 1gram precision std 11.54 12.83 12.64
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 108.85 112.65 112.27
Bleu ref len std 45.86 43.78 44.25
Bleu pred len mean 113.57 117.39 117.19
Bleu pred len std 47.70 45.75 46.51
Bleu bp mean 0.99 1.00 0.99
Bleu bp std 0.05 0.03 0.04
Bleu ratio mean 1.05 1.04 1.04
Bleu ratio std 0.09 0.07 0.08
Chrf score mean 9.91 10.46 10.32
Chrf score std 11.91 13.19 12.99
Ter score mean 100.59 99.24 99.44
Ter score std 18.73 17.23 17.20
Chrfpp score mean 9.91 10.46 10.32
Chrfpp score std 11.91 13.19 12.99
Table 52: BLEU, ChrF, ChrF++, and TER results for Japanese on the AI papers and model cards evaluation set using hf-seamless-m4t-large.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 10.55 10.60 10.58
Bleu score std 12.87 12.75 12.82
Bleu 1gram precision mean 10.58 10.63 10.61
Bleu 1gram precision std 12.99 12.85 12.91
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 116.51 119.64 120.03
Bleu ref len std 46.42 44.58 44.87
Bleu pred len mean 121.10 124.69 125.10
Bleu pred len std 48.28 46.61 46.94
Bleu bp mean 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.04 1.04 1.04
Bleu ratio std 0.05 0.05 0.05
Chrf score mean 10.99 11.05 11.03
Chrf score std 13.18 13.04 13.11
Ter score mean 98.39 98.32 98.27
Ter score std 18.10 17.19 17.31
Chrfpp score mean 10.99 11.05 11.03
Chrfpp score std 13.18 13.04 13.11
Table 53: BLEU, ChrF, ChrF++, and TER results for Japanese on the AI papers and model cards evaluation set using gpt-4o-mini.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 9.22 9.45 9.57
Bleu score std 12.23 12.32 12.84
Bleu 1gram precision mean 9.25 9.48 9.60
Bleu 1gram precision std 12.35 12.44 12.95
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 114.14 115.01 114.24
Bleu ref len std 46.74 46.44 46.64
Bleu pred len mean 120.26 121.02 120.36
Bleu pred len std 49.41 48.96 49.29
Bleu bp mean 1.00 1.00 1.00
Bleu bp std 0.02 0.02 0.02
Bleu ratio mean 1.06 1.06 1.06
Bleu ratio std 0.08 0.08 0.08
Chrf score mean 9.68 9.93 10.05
Chrf score std 12.48 12.60 13.09
Ter score mean 100.30 99.87 99.72
Ter score std 17.73 17.26 17.76
Chrfpp score mean 9.68 9.93 10.05
Chrfpp score std 12.48 12.60 13.09
Table 54: BLEU, ChrF, ChrF++, and TER results for Japanese on the AI papers and model cards evaluation set using aya-23-8B.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 8.11 8.08 8.16
Bleu score std 10.28 10.12 9.93
Bleu 1gram precision mean 8.12 8.09 8.17
Bleu 1gram precision std 10.29 10.13 9.94
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 119.56 119.59 119.77
Bleu ref len std 46.73 46.56 46.72
Bleu pred len mean 126.61 126.63 126.85
Bleu pred len std 48.80 48.64 48.80
Bleu bp mean 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.06 1.06 1.06
Bleu ratio std 0.07 0.07 0.07
Chrf score mean 8.59 8.57 8.65
Chrf score std 10.70 10.55 10.35
Ter score mean 101.99 101.96 101.92
Ter score std 16.22 15.81 15.79
Chrfpp score mean 8.59 8.57 8.65
Chrfpp score std 10.70 10.55 10.35
Table 55: BLEU, ChrF, ChrF++, and TER results for Japanese on the AI papers and model cards evaluation set using aya-expanse-8B.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 19.18 19.07 18.94
Bleu score std 19.82 19.23 19.06
Bleu 1gram precision mean 19.24 19.13 19.00
Bleu 1gram precision std 19.82 19.25 19.07
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 195.84 220.60 200.93
Bleu ref len std 83.60 84.78 84.41
Bleu pred len mean 196.94 221.81 202.15
Bleu pred len std 84.71 85.63 85.45
Bleu bp mean 0.99 1.00 1.00
Bleu bp std 0.03 0.02 0.02
Bleu ratio mean 1.00 1.00 1.00
Bleu ratio std 0.03 0.03 0.03
Chrf score mean 19.63 19.56 19.43
Chrf score std 19.96 19.44 19.28
Ter score mean 93.26 93.54 93.63
Ter score std 24.11 23.06 22.85
Chrfpp score mean 19.63 19.56 19.43
Chrfpp score std 19.96 19.44 19.28
Table 56: BLEU, ChrF, ChrF++, and TER results for Russian on the AI papers and model cards evaluation set using nllb-200-3.3B.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 18.43 19.38 18.78
Bleu score std 17.90 18.51 17.46
Bleu 1gram precision mean 18.48 19.44 18.84
Bleu 1gram precision std 17.92 18.54 17.49
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 222.40 231.72 229.59
Bleu ref len std 84.52 83.39 82.20
Bleu pred len mean 223.39 232.74 230.60
Bleu pred len std 85.50 84.23 83.10
Bleu bp mean 0.99 1.00 1.00
Bleu bp std 0.03 0.02 0.02
Bleu ratio mean 1.00 1.00 1.00
Bleu ratio std 0.03 0.03 0.03
Chrf score mean 18.88 19.86 19.26
Chrf score std 18.04 18.68 17.62
Ter score mean 94.15 93.12 93.77
Ter score std 22.06 22.29 21.14
Chrfpp score mean 18.88 19.86 19.26
Chrfpp score std 18.04 18.68 17.62
Table 57: BLEU, ChrF, ChrF++, and TER results for Russian on the AI papers and model cards evaluation set using hf-seamless-m4t-large.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 18.02 19.54 18.43
Bleu score std 19.69 20.97 19.57
Bleu 1gram precision mean 18.13 19.62 18.51
Bleu 1gram precision std 19.72 21.00 19.60
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 230.39 236.14 236.37
Bleu ref len std 90.66 85.39 85.34
Bleu pred len mean 230.15 236.19 236.33
Bleu pred len std 90.89 85.82 85.72
Bleu bp mean 0.99 0.99 0.99
Bleu bp std 0.04 0.03 0.03
Bleu ratio mean 1.00 1.00 1.00
Bleu ratio std 0.03 0.02 0.02
Chrf score mean 18.44 19.95 18.83
Chrf score std 19.83 21.10 19.70
Ter score mean 94.40 92.58 93.89
Ter score std 23.67 24.65 23.35
Chrfpp score mean 18.44 19.95 18.83
Chrfpp score std 19.83 21.10 19.70
Table 58: BLEU, ChrF, ChrF++, and TER results for Russian on the AI papers and model cards evaluation set using gpt-4o-mini.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 16.39 17.76 17.01
Bleu score std 17.61 18.41 18.37
Bleu 1gram precision mean 16.47 17.84 17.09
Bleu 1gram precision std 17.69 18.48 18.45
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 232.81 234.38 233.02
Bleu ref len std 88.09 86.95 88.03
Bleu pred len mean 232.90 234.61 233.12
Bleu pred len std 88.38 87.21 88.27
Bleu bp mean 0.99 0.99 0.99
Bleu bp std 0.02 0.02 0.02
Bleu ratio mean 1.00 1.00 1.00
Bleu ratio std 0.03 0.03 0.03
Chrf score mean 16.78 18.16 17.40
Chrf score std 17.75 18.54 18.50
Ter score mean 96.28 94.55 95.44
Ter score std 21.39 21.92 21.93
Chrfpp score mean 16.78 18.16 17.40
Chrfpp score std 17.75 18.54 18.50
Table 59: BLEU, ChrF, ChrF++, and TER results for Russian on the AI papers and model cards evaluation set using aya-23-8B.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 13.84 14.06 14.17
Bleu score std 13.57 14.08 13.26
Bleu 1gram precision mean 13.89 14.10 14.21
Bleu 1gram precision std 13.61 14.12 13.31
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 238.17 238.01 238.42
Bleu ref len std 87.01 87.13 87.09
Bleu pred len mean 240.60 240.41 240.79
Bleu pred len std 86.86 87.01 86.90
Bleu bp mean 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.01 1.01 1.01
Bleu ratio std 0.04 0.04 0.04
Chrf score mean 14.32 14.53 14.66
Chrf score std 13.83 14.31 13.54
Ter score mean 99.77 99.58 99.37
Ter score std 17.18 17.63 16.80
Chrfpp score mean 14.32 14.53 14.66
Chrfpp score std 13.83 14.31 13.54
Table 60: BLEU, ChrF, ChrF++, and TER results for Russian on the AI papers and model cards evaluation set using aya-expanse-8B.
Direct Translation Prompting Refinement Word Alignment
nllb-200-3.3B
Arabic 0.81 0.83 0.81
Chinese 0.80 0.82 0.81
French 0.79 0.83 0.77
Japanese 0.83 0.85 0.84
Russian 0.80 0.83 0.78
hf-seamless-m4t-large
Arabic 0.82 0.83 0.82
Chinese 0.81 0.83 0.81
French 0.84 0.85 0.83
Japanese 0.84 0.85 0.84
Russian 0.83 0.85 0.81
gpt-4o-mini
Arabic 0.84 0.84 0.83
Chinese 0.86 0.86 0.86
French 0.85 0.85 0.84
Japanese 0.88 0.88 0.88
Russian 0.85 0.85 0.83
aya-23-8B
Arabic 0.82 0.83 0.81
Chinese 0.84 0.85 0.84
French 0.83 0.84 0.82
Japanese 0.84 0.85 0.84
Russian 0.83 0.84 0.81
aya-expanse-8B
Arabic 0.81 0.81 0.80
Chinese 0.82 0.83 0.82
French 0.81 0.82 0.80
Japanese 0.85 0.86 0.86
Russian 0.83 0.83 0.81
Table 61: COMET results on the AI papers and model cards evaluation set.